Laser Subcommittee Meeting 2018-08-27


  • Blake, NickB, Fletcher


  • Little red
    • Bed out of level.
    • Reports of z-axis belt slip over last month. Unsure if z-probe occasionally missing and head crashing or just tensioners slipping.
    • Removed and thoroughly cleaned the tensioners and surface they are attached to. Cleaned fasteners.
    • Relevelled bed
    • Firmly tightened bolts for tensioners to prevent slipping
    • Checked heights and adjusted offset from 9mm to 10mm and which gave best cut (head height 18mm)
    • Cutting at 13mm/s for 3mm MDF
    • Still overcurrent at 24mA
    • Still tube change being pursued
      • Require end extension
      • Require tube mounts
  • Big red
    • Hour meter 295.5 hrs
    • Relevelled the bed
    • Adjusted offset 16mm. Nozzle to target height 21mm.
    • Cleaned lens (was a bit misty)
  • Kellogg the KH7050B
    • Door sensors
      • Further investigation over last month showed the door sensors do not work now as they were thought to. They are no longer emitting a 10kHz square wave. Various tests were performed several weeks ago with Steve, Blake and Kamil.
      • Steve tried to find more information on sensors but was unable to
      • Probably replace with magnet door sensors
    • Discussion on z-axis drive
      • Kamil's large stepper arrived but was decided its too likely to strip belt until pole gears are replace with larger ones
    • Need to get a spare set of belts for Kellogg too
  • subcommittee/laser-minutes-20180827.txt
  • Last modified: 2020/03/12 12:49
  • (external edit)